### v1

@@ Abstract
Identifying vulnerabilities in source code is crucial, especially in critical software components. Existing methods such as static analysis, dynamic analysis, formal verification, and recently Large Language Models are widely used to detect security flaws. This paper introduces CASTLE (CWE Automated Security Testing and Low-Level Evaluation), a benchmarking framework for evaluating the vulnerability detection capabilities of different methods. We assess 13 static analysis tools, 10 LLMs, and 2 formal verification tools using a hand-crafted dataset of 250 micro-benchmark programs covering 25 common CWEs. We propose the CASTLE Score, a novel evaluation metric to ensure fair comparison. Our results reveal key differences: ESBMC (a formal verification tool) minimizes false positives but struggles with vulnerabilities beyond model checking, such as weak cryptography or SQL injection. Static analyzers suffer from high false positives, increasing manual validation efforts for developers. LLMs perform exceptionally well in the CASTLE dataset when identifying vulnerabilities in small code snippets. However, their accuracy declines, and hallucinations increase as the code size grows. These results suggest that LLMs could play a pivotal role in future security solutions, particularly within code completion frameworks, where they can provide real-time guidance to prevent vulnerabilities. The dataset is accessible at https://github.com/CASTLE-Benchmark.
@@ 1 Introduction
Rapid advancements in artificial intelligence (AI) have sparked both excitement and concern about the future of traditional software engineering. For instance, Meta’s recent announcement that AI could soon replace many software engineering roles highlights a shifting landscape in code development [1]. While AI-driven code generation offers remarkable efficiency, a study by Tihanyi et al. found that all examined large language models (LLMs) produced vulnerable C code [2]. Similar conclusions have been reached in studies examining other programming languages, such as PHP and Python [3, 4]. These large-scale studies consistently indicate that such vulnerabilities arise, in part, because LLMs lack the contextual understanding during the generation process. Several studies highlight that once the code is generated, and a vulnerability is identified, LLMs excel in fixing these [5, 6]. The question is, how do we find the vulnerabilities? Despite the growing importance of automated software verification, developers and security practitioners lack clear guidance on which tools are most reliable for detecting vulnerabilities in C code. Several interrelated issues contribute to this uncertainty:
@@
Given these challenges, a robust and compilable benchmark dataset that accurately captures major CWE vulnerabilities is paramount in order to answer the following research questions:
RQ1: How do state-of-the-art static analysis tools, formal verification methods, and LLM-based approaches compare in effectively detecting vulnerabilities in C code? RQ2: Are combinations of tools more effective than using a single tool? RQ3: What metrics can reliably demonstrate these differences among various tools?
@@ 2 Related work
Ensuring software correctness, safety, and security is central to software engineering. Examining related literature on the role of AI in software development, most of the existing work and benchmarking approaches focused on testing LLMs capabilities in producing functionally correct code. However, safety and security is just as important.
@@ 2.1 Datasets and Benchmarks
Furthermore, a key challenge is that many popular datasets lack compilable programs, making it impossible to meaningfully assess formal verification tools such as the Efficient SMT-based Context-Bounded Model Checker (ESBMC). Among compilable options, SARD stands out for its inclusion of the Juliet test cases and 45,437 C samples mapped to CWE categories. However, some files exceed 3,000 lines of code, creating two constraints:
@@
In contrast, CASTLE provides a collection of compilable code snippets, deliberately crafted to cover major CWEs while minimizing the number of queries required for effective analysis. This design enables straightforward deployment of both LLM-based methods and traditional static analyzers with specialized wrappers, facilitating rapid, automated evaluation across a variety of tools. Additionally, the newly introduced CASTLE score provides a more detailed comparative metric than conventional pass/fail assessments, allowing for clearer differentiation of subtle performance variations among state-of-the-art tools. The CASTLE dataset balances both vulnerable and non-vulnerable samples, permitting more robust evaluations of false positives and false negatives alike.
@@ 2.3 LLM-Based Vulnerability Detection
Recent years have witnessed a growing interest in using Large Language Models (LLMs) for automated vulnerability detection [25, 26, 27]. Although these models are often praised for their ability to handle diverse code repositories, they primarily rely on pattern-based sequence learning, rather than neuro-symbolic reasoning. As a result, LLMs can detect certain coding flaws effectively, yet they remain susceptible to overlooking complex or context-dependent vulnerabilities.
@@
Vulnerability detectors typically leverage transformer-based code models trained on massive code corpora spanning multiple languages. These training datasets often contain their own share of insecure code, potentially introducing biases or “model collapse” issues [28]. Broadly, transformer models are categorized into three groups [29]:
@@ 3 Methodology
Figure 1: The CASTLE Benchmark Framework
This section overviews the dataset creation process and introduces the newly developed evaluation metrics used in our research. Figure 1 provides a visual overview of the dataset creation and testing framework.
@@ 3.1 Dataset
The CASTLE dataset comprises 250 small programs in C, each crafted manually by cybersecurity experts. It encompasses 25 distinct CWEs, with exactly 10 test cases per CWE (6 vulnerable and 4 non-vulnerable). This balanced distribution facilitates focused assessments of each tool’s vulnerability detection capabilities while accurately measuring false positives. In ambiguous cases, experts selected a higher-level CWE category or iteratively refined the test until only the most relevant CWE remained.
Each program was required to compile without errors—although compiler warnings were permitted—as this is a requirement for formal verification tools. All benchmarks were written in C, selected for its capacity to accommodate a wide range of vulnerability types, including intricate memory management issues. Furthermore, each test case was restricted to a single file (with optional external libraries) and designed to contain exactly one or zero vulnerabilities. This structure simplifies the identification of vulnerabilities and helps prevent confusion when validating false positives.
@@
Table 2: List of CWEs in the benchmark in relation to their presence in MITRE’s 2024 Top 25 vulnerability list [30]
When incorporating the system prompt alongside the source code, the total input tokens across the dataset amount to approximately 115,620 tokens using the cl100k_base encoding scheme. This total reflects the resource considerations required when running evaluations with token-sensitive language models. The dataset was intentionally capped at 250 benchmarks to make thorough manual verification feasible. This rigorous verification process is indispensable for detecting false positives and confirming line-level detections. Moreover, this selective approach supports cost-effective evaluation of computationally intensive tools, including advanced LLMs (e.g., GPT-o1, GPT-o3, DeepSeek R1).
The benchmarks exhibit substantial variability in complexity. Code lengths range from 7 to 164 lines, yielding a total of 10,392 lines (an average of 42 lines). Each includes 1–8 functions (2.2 on average), with cyclomatic complexity values spanning 1–29 (mean 6.3). Halstead volumes range from approximately 89.9 to over 5,246.7, averaging 1,104.8.
This breadth ensures the dataset covers a wide spectrum of vulnerabilities, from lower-level issues (e.g., memory management flaws, race conditions) to higher-level security risks (e.g., command injections, hard-coded credentials). Most CWEs were chosen based on their prevalence in the Top 25 CWEs of 2023–2024. To ensure overall quality and reliability, each test underwent iterative validation by human experts. Table 2 provides a comprehensive list of the included CWEs.
@@ 3.2 Test Format and Wrappers
Additionally, the metadata specifies the CWE classification of the vulnerability, along with other contextual information. After processing, each test is converted into a JSON-formatted dictionary that includes the code, metadata, and computed software metrics (e.g., cyclomatic complexity, Halstead volume). This unified structure simplifies integration with the various wrappers, facilitating automated execution and standardized result reporting.
@@
[…]
To ensure a uniform and reproducible evaluation across all tools, we developed custom wrappers that automate installation, configuration, execution, and result retrieval. For freely available static analyzers (e.g., Clang Analyzer, ESBMC), each tool was containerized via Docker alongside its dependencies. We then employed Python scripts to run each tool on all the test cases, collecting and parsing results into a standardized report format.
@@
LLM-based vulnerability detection was driven by a generic script that interacted with standard OpenAI APIs. Each model was prompted to return JSON-formatted detection results. Smaller models (fewer than 6B parameters) often struggled to generate well-structured JSON, suggesting limitations in handling detailed output formats. Additionally, models were sensitive to line-specific detections, occasionally identifying the correct vulnerability but offsetting the line number. To address minor positioning errors, we also prompted LLMs to provide the corresponding code lines, allowing minimal adjustments during evaluation.
@@
[!htbp] rompt Template for LLMs for JSON-formatted vulnerability reports
LLM System Prompt 1You are a professional security analyst reviewing C code for vulnerabilities. 2You will list the found vulnerabilities in a JSON format using the exact template below: 3 4‘‘‘ 5[ 6 {"severity": string, "line": int, "cwe": int, "message": string, "line_content": string}, 7 ... 8] 9‘‘‘ 10### Rules: 111. Do not omit the triple backticks (‘‘‘ at the beginning and ‘‘‘ at the end). 122. If you did not find any vulnerabilities, return an empty list. 133. If you dont know the CWE number, set it to 0. 144. Any response that does not follow the above format is invalid. 155. You get 5 points for each vulnerability you find, but get -1 for all false positives you report. 16 17Now, review the following C code and return your response:
@@ 3.3 CASTLE Score
In this section, we introduce the CASTLE score, a new metric for evaluating the performance of vulnerability detection tools with the CASTLE-Benchmark. The CASTLE score integrates both true- and false-positive rates, awards bonus points for detecting high-impact vulnerabilities (based on the Top 25 CWEs), and rewards correct identification of non-vulnerable code. By incorporating these factors, the metric more finely captures a tool’s overall reliability than standard pass/fail evaluations.
@@
  • All Other Cases. If the tool misses a vulnerability (failing to report $v_{i}$), or incorrectly flags any vulnerability (including false positives in a non-vulnerable test), the score is reduced by one for each false-positive finding ($-\lvert t(d_{i})\rvert$). Notably, if the tool reports nothing on a vulnerable benchmark, it does not incur any additional penalties.
@@ 4 Discussion
The CASTLE Score is designed to provide a balanced assessment of a tool’s effectiveness by considering both true and false positives, as well as the severity of vulnerabilities. Consequently, not finding a high-severity vulnerability leads to larger penalties compared to less impactful ones. A negative CASTLE Score could indicate that the volume of false positives generated by a tool imposes a significant triage burden on developers, outweighing its potential benefits. Overall, both the benchmark dataset and the introduced evaluation metric helped to highlight the strengths and weaknesses of various static analyzers.
@@ 4.1 Tool evaluation on CASTLE Benchmark
Figure 2: CASTLE Scores for tools tested on 250 C programs, including the top five tool combinations. Tools reporting no issues score 200 points. The theoretical maximum of perfect score is 1250 points.
@@
The highest-performing individual tool in our analysis was ESBMC, a formal verification tool. Formal Verification methods have a main disadvantage of not being able to detect non-formal issues, such as SQL Injection, Path traversal or hard-coded credentials, however they compensate for this with their low false positive rate. Theoretically, bounded model checkers cannot produce false positives, as they always provide a counterexample to their findings, except in cases where the tool itself has bugs or reports esoteric scenarios. CBMC achieved 0 false positives, the only tool to do so (see Figure 3). ESBMC, however, had 32 false positives in three categories, some of which we submitted as bug reports to the project [31] [32] [33] [34]. While this dataset with is short code samples allowed relaxed setting for ESBMC with a longer timeout; --overflow --no-unwinding-assertions --memory-leak-check --timeout 60 --multi-property --show-stacktrace, with larger codebases formal verification tools could peotentially struggle to finish the verification process in reasonalbe time, thereby limiting their thoroughness and reliability in giving accurate results. The best performing SAST tool is CodeQL, which found 29% of the weaknesses in the code (45/150), which is the highest of the average of 23% among other SASTs. SonarQube also received 29%, but it was dragged down by reporting 2.7 times as many false positives than CodeQL. Several tools, including Clang Analyzer and CBMC, displayed extremely high precision (87% and 100% respectively) but struggled with low recall (9% and 12%). This trade-off implies they excel at correctly labeling the few issues they do detect, yet they fail to identify a substantial portion of vulnerabilities. Conversely, ESBMC’s more balanced approach (62% precision, 35% recall) often provides a more reliable day-to-day detection rate for developers. Tools such as Splint and CodeThreat generated exceptionally high numbers of false positives (1,027 and 1,101, respectively). Their resulting negative CASTLE Scores (-598 and -692) illustrate how overwhelming false alerts can erode a tool’s practical utility. Although both tools still produced a modest number of true positives, the excessive manual triage effort likely outweighs any marginal benefits for most real-world projects.
@@
Another advantage of the CASTLE score, over traditional metrics, is that it provides a comparison between using tool combinations. If a pair of tools has a high overlap in the CWEs they can detect, the CASTLE Score of their combination will yield a lower result than the individual tools because of the oversized impact of increasing the rate of false positives. When looking at combination scores, the biggest increase happens when ESBMC and CodeQL is combined, yielding 814 points. This is a 153 point increase over the higher performing ESBMC’s base score of 661, and a 23% increase in the effectiveness of using both tools instead of just ESBMC, with a 28% increase above just using CodeQL. This shows that selecting tool combinations with different strengths significantly boosts the efficacy of the static analysis process.

### v2

@@ Abstract
Identifying vulnerabilities in source code is crucial, especially in critical software components. Existing methods such as static analysis, dynamic analysis, formal verification, and recently Large Language Models are widely used to detect security flaws. This paper introduces CASTLE (CWE Automated Security Testing and Low-Level Evaluation), a benchmarking framework for evaluating the vulnerability detection capabilities of different methods. We assess 13 static analysis tools, 10 LLMs, and 2 formal verification tools using a hand-crafted dataset of 250 micro-benchmark programs covering 25 common CWEs. We propose the CASTLE Score, a novel evaluation metric for fair comparison. Our results reveal key differences: ESBMC (a formal verification tool) minimizes false positives but struggles with vulnerabilities beyond model checking, such as weak cryptography or SQL injection. Static analyzers suffer from high false positives, increasing manual validation efforts for developers. LLMs perform exceptionally well in the CASTLE dataset when identifying vulnerabilities in small code snippets. However, their accuracy declines, and hallucinations increase as the code size grows. These results suggest that LLMs could play a pivotal role in future security solutions, particularly within code completion frameworks, where they can provide real-time guidance to prevent vulnerabilities. The dataset is accessible at https://github.com/CASTLE-Benchmark .
@@ 1 Introduction
Rapid advancements in artificial intelligence (AI) have sparked both excitement and concern about the future of traditional software engineering. For instance, Meta’s recent announcement that AI could soon replace many software engineering roles highlights a shifting landscape in code development [1]. While AI-driven code generation offers remarkable efficiency, a study by Tihanyi et al. found that all examined large language models (LLMs) produced vulnerable C code [2]. Similar conclusions have been reached in studies examining other programming languages, such as PHP and Python [3, 4]. These large-scale studies consistently indicate that such vulnerabilities arise partly because LLMs lack contextual understanding during the generation process. Several studies highlight that once the code is generated, and a vulnerability is identified, LLMs excel in fixing these [5, 6]. The question is, how do we find the vulnerabilities? Despite the growing importance of automated software verification, developers and security practitioners lack clear guidance on which tools are most reliable for detecting vulnerabilities in C code. Several interrelated issues contribute to this uncertainty:
@@
Given these challenges, a robust and compilable benchmark dataset that accurately captures major CWE vulnerabilities is paramount to answer the following research questions:
RQ1: How do state-of-the-art static analysis tools, formal verification methods, and LLM-based approaches compare to effectively detecting C code vulnerabilities? RQ2: Are combinations of tools more effective than using a single tool? RQ3: What metrics can reliably demonstrate these differences among various tools?
@@ 2 Related work
Ensuring software correctness, safety, and security is central to software engineering. Examining related literature on the role of AI in software development, most of the existing work and benchmarking approaches focused on testing LLMs’ capabilities in producing functionally correct code. However, safety and security are just as important.
@@ 2.1 Datasets and Benchmarks
Furthermore, a key challenge is that many popular datasets lack compilable programs, making it impossible to meaningfully assess formal verification tools such as the Efficient SMT-based Context-Bounded Model Checker (ESBMC). Among compilable options, SARD stands out for including the Juliet test cases and $45,437$ C samples mapped to CWE categories. However, some files exceed $3,000$ lines of code, creating two constraints:
@@
In contrast, CASTLE provides a collection of compilable code snippets, deliberately crafted to cover major CWEs while minimizing the number of queries required for effective analysis. This design enables the straightforward deployment of LLM-based methods and traditional static analyzers with specialized wrappers, facilitating rapid, automated evaluation across various tools. Additionally, the newly introduced CASTLE score provides a more detailed comparative metric than conventional pass/fail assessments, allowing for clearer differentiation of subtle performance variations among state-of-the-art tools. The CASTLE dataset balances vulnerable and non-vulnerable samples, permitting more robust evaluations of false positives and negatives.
@@ 2.3 LLM-Based Vulnerability Detection
Recent years have witnessed a growing interest in using Large Language Models (LLMs) for automated vulnerability detection [25, 26, 27]. Although these models are often praised for handling diverse code repositories, they primarily rely on pattern-based sequence learning rather than (neuro-)symbolic reasoning. As a result, LLMs can detect certain coding flaws effectively, yet they remain susceptible to overlooking complex or context-dependent vulnerabilities.
@@
Vulnerability detectors typically leverage transformer-based code models trained on massive code corpora spanning multiple languages. These training datasets often contain their share of insecure code, potentially introducing biases or “model collapse” issues [28]. Broadly, transformer models are categorized into three groups [29]:
@@ 3 Methodology
Figure 1: The CASTLE Benchmark Framework.
This section overviews the dataset creation process and introduces our research’s newly developed evaluation metrics. Figure 1 provides a visual overview of the dataset creation and testing framework.
@@ 3.1 Dataset
The CASTLE dataset comprises 250 small programs in C, each crafted manually by cybersecurity experts. It encompasses $25$ distinct CWEs, with $10$ test cases per CWE ($6$ vulnerable and $4$ non-vulnerable). This balanced distribution facilitates focused assessments of each tool’s vulnerability detection capabilities while accurately measuring false positives. In ambiguous cases, experts selected a higher-level CWE category or iteratively refined the test until only the most relevant CWE remained.
Each program was required to compile without errors—although compiler warnings were permitted—as this is a requirement for formal verification tools. All benchmarks were written in C and selected for their capacity to accommodate a wide range of vulnerability types, including intricate memory management issues. Furthermore, each test case was restricted to a single file (with optional external libraries) and designed to contain exactly one or zero vulnerabilities. This structure simplifies the identification of vulnerabilities and helps prevent confusion when validating false positives.
@@
Table 2: List of CWEs in the benchmark concerning their presence in MITRE’s 2024 Top 25 vulnerability list [30]
When incorporating the system prompt alongside the source code, the total input tokens across the dataset amount to approximately 115,620 tokens using the cl100k_base encoding scheme. This total reflects the resource considerations required when running evaluations with token-sensitive language models. The dataset was intentionally capped at 250 benchmarks to make thorough manual verification feasible. This rigorous verification process is indispensable for detecting false positives and confirming line-level detections. Moreover, this selective approach supports the cost-effective evaluation of computationally intensive tools, including advanced LLMs (e.g., GPT-o1, GPT-o3, DeepSeek R1).
The benchmarks exhibit substantial variability in complexity. Code lengths range from $7$ to $164$ lines, yielding $10,392$ lines (an average of $42$ lines). Each includes 11118888 functions ($2.2$ on average), with cyclomatic complexity values spanning 1–29 (mean $6.3$). Halstead volumes range from approximately $89.9$ to over $5,246.7$, averaging $1,104.8$.
This breadth ensures the dataset covers a wide spectrum of vulnerabilities, from lower-level issues (e.g., memory management flaws, race conditions) to higher-level security risks (e.g., command injections, hard-coded credentials). Most CWEs were chosen based on their prevalence in the Top 25 CWEs of 2023–2024. Each test underwent iterative validation by human experts to ensure overall quality and reliability. Table 2 provides a comprehensive list of the included CWEs.
@@ 3.2 Test Format and Wrappers
Additionally, the metadata specifies the vulnerability’s CWE classification and other contextual information. After processing, each test is converted into a JSON-formatted dictionary that includes the code, metadata, and computed software metrics (e.g., cyclomatic complexity, Halstead volume). This unified structure simplifies integration with the various wrappers, facilitating automated execution and standardized result reporting.
@@
CASTLE-787-1.c Test Source Code {minted} [xleftmargin=6pt,linenos,escapeinside=||]c /* ================================================ dataset: CASTLE-Benchmark name: CASTLE-787-1.c version: 1.1 compile: gcc CASTLE-787-1.c -o CASTLE-787-1 vulnerable: true description: Buffer overflow in scanf function copying into a fixed length buffer. cwe: 787 ============================================== */ #include <stdio.h> int main(int argc, char *argv[]) char reg_name[12]; printf("Enter your username:"); scanf("printf("Hello return 0;
To ensure a uniform and reproducible evaluation across all tools, we developed custom wrappers that automate installation, configuration, execution, and result retrieval. Each tool was containerized via Docker alongside its dependencies for freely available static analyzers (e.g., Clang Analyzer, ESBMC). We then used Python scripts to run each tool on all test cases, collecting and parsing the results into a standardized report format.
@@
LLM-based vulnerability detection was driven by a generic script that interacted with standard OpenAI APIs. Each model was prompted to return JSON-formatted detection results. Smaller models (fewer than 6B parameters) often struggled to generate well-structured JSON, suggesting limitations in handling detailed output formats. Additionally, models were sensitive to line-specific detections, occasionally identifying the correct vulnerability but offsetting the line number. We also prompted LLMs to provide the corresponding code lines to address minor positioning errors, allowing minimal adjustments during evaluation.
@@
[!htbp] Prompt Template for LLMs for JSON-formatted vulnerability reports
LLM System Prompt {minted} [xleftmargin=6pt,linenos,breaklines,escapeinside=||]md You are a professional security analyst reviewing C code for vulnerabilities. You will list the found vulnerabilities in a JSON format using the exact template below: “‘ [ "severity": string, "line": int, "cwe": int, "message": string, "line_content": string, … ] “‘ ### Rules: 1. Do not omit the triple backticks (“‘ at the beginning and “‘ at the end). 2. If you did not find any vulnerabilities, return an empty list. 3. If you don’t know the CWE number, set it to 0. 4. Any response that does not follow the above format is invalid. 5. You get 5 points for each vulnerability you find, but get -1 for all false positives you report. Now, review the following C code and return your response:
@@ 3.3 CASTLE Score
In this section, we introduce the CASTLE score, a new metric for evaluating the performance of vulnerability detection tools with the CASTLE-Benchmark. The CASTLE score integrates both true- and false-positive rates, awards bonus points for detecting high-impact vulnerabilities (based on the Top 25 CWEs), and rewards correct identification of non-vulnerable code. By incorporating these factors, the metric better captures a tool’s overall reliability than standard pass/fail evaluations.
@@
  • All Other Cases. If the tool misses a vulnerability (failing to report $v_{i}$), or incorrectly flags any vulnerability (including false positives in a non-vulnerable test), the score is reduced by one for each false-positive finding ($-\lvert t(d_{i})\rvert$). Notably, it does not incur additional penalties if the tool reports nothing on a vulnerable benchmark.
@@ 4 Discussion
The CASTLE Score is designed to provide a balanced assessment of a tool’s effectiveness by considering both true and false positives and the severity of vulnerabilities. Consequently, not finding a high-severity vulnerability leads to larger penalties than less impactful ones. A negative CASTLE Score could indicate that the volume of false positives generated by a tool imposes a significant triage burden on developers, outweighing its potential benefits. Overall, both the benchmark dataset and the introduced evaluation metric helped highlight various static analyzers’ strengths and weaknesses.
@@ 4.1 Tool evaluation on CASTLE Benchmark
Figure 2: CASTLE Scores for tools tested on $250$ C programs, including the top five tool combinations. Tools reporting no issues score $200$ points. The theoretical maximum of a perfect score is $1250$ points.
@@
The highest-performing individual tool in our analysis was ESBMC, a formal verification tool. Formal Verification methods have the main disadvantage of being unable to detect non-formal issues, such as SQL Injection, Path traversal, or hard-coded credentials. However, they compensate for this with their low false positive rate. Theoretically, bounded model checkers cannot produce false positives, as they always provide a counterexample to their findings, except in cases where the tool itself has bugs or reports esoteric scenarios. CBMC achieved 00 false positives, the only tool to do so (see Figure 3). ESBMC, however, had $32$ false positives in three categories, some of which we submitted as bug reports to the project [31] [32] [33] [34]. While this dataset with is short code samples allowed relaxed setting for ESBMC with a longer timeout; --overflow --no-unwinding-assertions --memory-leak-check --timeout 60 --multi-property --show-stacktrace, with larger codebases formal verification tools could peotentially struggle to finish the verification process in reasonalbe time, thereby limiting their thoroughness and reliability in giving accurate results. The best-performing SAST tool is CodeQL, which found 29% of the weaknesses in the code (45/150), the highest of the average of 23% among other SASTs. SonarQube also received 29%, but it was dragged down by reporting 2.7 times as many false positives than CodeQL. Several tools, including Clang Analyzer and CBMC, displayed extremely high precision (87% and 100% respectively) but struggled with low recall (9% and 12%). This trade-off implies they excel at correctly labeling the few issues they detect, yet they fail to identify a substantial portion of vulnerabilities. Conversely, ESBMC’s more balanced approach (62% precision, 35% recall) often provides a more reliable day-to-day detection rate for developers. Splint and CodeThreat generated exceptionally high false positives (1,027 and 1,101, respectively). Their negative CASTLE Scores (-598 and -692) illustrate how overwhelmingly false alerts can erode a tool’s practical utility. Although both tools still produced a modest number of true positives, the excessive manual triage effort likely outweighs any marginal benefits for most real-world projects.
@@
Another advantage of the CASTLE score over traditional metrics is that it provides a comparison between using tool combinations. If a pair of tools has a high overlap in the CWEs they can detect, the CASTLE Score of their combination will yield a lower result than the individual tools because of the oversized impact of increasing the rate of false positives. When looking at combination scores, the biggest increase happens when ESBMC and CodeQL is combined, yielding 814 points. This is a 153 point increase over the higher performing ESBMC’s base score of 661, and a 23% increase in the effectiveness of using both tools instead of just ESBMC, with a 28% increase above just using CodeQL. This shows that selecting tool combinations with different strengths significantly boosts the efficacy of the static analysis process.